Suppose you invest money in two accounts. One of the accounts pay 8% annual interest, whereas the other pays 9% annual interest. If you have $8,000 more invested at 9% than you invested at 8%, how much do you have invested in each account if the total amount of interest you earn in a year is $1,060?

Answer by mananth(16946)

One of the accounts pay 8% annual interest,-----------x
other pays 9% -----x+8000
the total amount of interest you earn in a year is $1,060
8%x+9%(x+8000) = 1060
multiply by 100
8x+9(x+8000( = 106000
8x+9x+72000=106000
17x=106000-72000
17x=34000
x=2000
2000+8000= 10,000
2000 at 8 %
10,000 at 9%
check
8%*2000 + 9%*10000=1060